# And yet it moves: why irrotational waves induce a mean Lagrangian drift

### v2

@@ Appendix A Rotational Waves
Adding an arbitrary vorticity to waves in the Lagrangian frame is relatively straightforward as vorticity is conserved on particles in two-dimensional inviscid flow. Thus any Lagrangian formulation presented above which holds for a various collection of particles can be easily modified to account for vortical waves. As an example, we direct the reader to (Pizzo et al., 2023, eq. 2.9) to see how the drift is kinematically constrained by an arbitrary vorticity. For convenience we rewrite it here
@@
using the definition of $\bm{A}$ and $q$. Since the $\alpha$ derivative is over a wavelength, we can convert this to a phase average, resulting in
$\Gamma=\oint\bm{A}\cdot{\rm d}\bm{\alpha}=\int_{\alpha}^{\alpha+\lambda}\dot{x}x_{\alpha}+\dot{y}y_{\alpha}\,{\rm d}\alpha=-\int_{-\infty}^{\beta}\langle\mathcal{J}q\rangle\,{\rm d}\beta^{\prime}\,,$ (70)
@@
where $U(\beta)$ and $\langle T\rangle$ are defined the same as before. As a quick check consider the Gerstner wave, which is exactly described by the circular trajectories and pressure
$x(\alpha,\beta,t)$ $=\alpha-Ae^{k\beta}\sin(k(\alpha-ct))\,,$ (72)
$y(\alpha,\beta,t)$ $=\beta+Ae^{k\beta}\cos(k(\alpha-ct))+\frac{1}{2}A^{2}k\,,$ (73)
$p(\beta,t)$ $=-g\beta+\frac{1}{2}A^{2}k^{2}c^{2}(e^{2k\beta}-1)\,,$ (74)
@@
which validates the result for this special case. Thus (71) is the general version of (25), and states that in addition to the mean Lagrangian drift, there can be a balance between the mean kinetic energy density and the vorticity of particles for periodic waves.
@@
Once again, we can integrate this equation and use the same argument to constrain the constant of integration $(f(\alpha,t)=c^{2}/2)$ to find

### v3

@@ Appendix A Rotational Waves
Due to the relative simplicity of irrotational planar flow in the Eulerian frame, allowing for the use of a velocity potential, the vast majority of surface wave literature concerns irrotational waves. However it is generally understood that common effects such as wave breaking (Rapp & Melville, 1990; Pizzo & Melville, 2013) and ocean-atmosphere shear flows which can generate surface waves (Young & Wolfe, 2014) imply that real world surface gravity waves often contain vorticity. Without the benefit of potential flow, Eulerian treatments of rotational surface gravity waves (Phillips, 2001; Constantin et al., 2006) become quite difficult. This is where the Lagrangian machinery outlined above operates best as vorticity is conserved on particles in two-dimensional inviscid flow. Thus any Lagrangian formulation presented above which holds for a various collection of particles can be easily modified to account for vortical waves. As an example, we direct the reader to (Pizzo et al., 2023, eq. 2.9) to see how the drift is dynamically constrained by an arbitrary vorticity. For convenience we rewrite it here
@@
using the definition of $\bm{A}$ and $q$. This is also equivalent to the enclosed vorticity (with a minus sign for the orientation of $\bm{\hat{n}}$) within the material loop, identical to that in the Eulerian frame. Since the $\alpha$ derivative is over a wavelength, we can convert this to a phase average, resulting in
$\Gamma=\oint\bm{A}\cdot{\rm d}\bm{\alpha}=\int_{\alpha}^{\alpha+\lambda}\dot{x}x_{\alpha}+\dot{y}y_{\alpha}\,{\rm d}\alpha^{\prime}=-\int_{-\infty}^{\beta}\langle\mathcal{J}q\rangle\,{\rm d}\beta^{\prime}\,,$ (70)
@@
where $U(\beta)$ and $\langle T\rangle$ are defined the same as before. As a quick check consider the Gerstner (1802) wave, which is exactly described by the circular trajectories and pressure
$x(\alpha,\beta,\tau)$ $=\alpha-Ae^{k\beta}\sin(k(\alpha-c\tau))\,,$ (72)
$y(\alpha,\beta,\tau)$ $=\beta+Ae^{k\beta}\cos(k(\alpha-c\tau))+\frac{1}{2}A^{2}k\,,$ (73)
$p(\beta,\tau)$ $=-g\beta+\frac{1}{2}A^{2}k^{2}c^{2}(e^{2k\beta}-1)\,,$ (74)
@@
which validates the result for this special case. Thus (71) is the generalization of (27), and states that there is actually a balance between drift, kinetic energy density, and vorticity for the waves considered. Stokes waves, where the balance is entirely between drift and kinetic energy density, or Gerstner waves, which balance kinetic energy density and vorticity, are thus limiting cases for this general result.
@@
Once again, we can integrate this equation and use the same argument to constrain the constant of integration $(f(\alpha,\tau)=c^{2}/2)$ to find